### v1

@@ 1 Introduction
Covariance matrix estimation is central in modern finance. It influences portfolio construction, risk management, and asset pricing models. Markowitz, (1952) introduced mean-variance optimisation, highlighting how correlation among assets shapes the risk-return trade-off. Beyond portfolio choice, covariance estimation underpins factor-based models, hedging strategies, and even debates on market efficiency, where accurate correlation forecasts can drive risk control and potential outperformance.
@@
However, GARCH-type approaches struggle with large cross-sections, falling prey to the curse of dimensionality (Robert F. Engle and Wolf,, 2019). Tools like principal component analysis and random matrix theory (Ledoit and Wolf,, 2003; Laloux et al.,, 2000) have been effective for large datasets. Recent advances, such as dynamic conditional correlation with nonlinear shrinkage (Robert F. Engle and Wolf,, 2019), perform well overall, as per De Nard et al., (2024). Still, most of these methods focus on very short-term horizons, often just one step ahead or one day ahead, and rely on high-frequency data.
@@
Long-horizon forecasts matter for institutional investors like pension funds and endowments, which rebalance portfolios over weeks or months. Moreover, high-frequency data are not always public or cost-effective, making the short-horizon focus less relevant for these institutions. Furthermore, slow-moving trends, such as equity–bond correlation shifts, can be overlooked if one relies solely on high-frequency econometric models. Sandoval and Franca, (2012) highlights how regime shifts or significant structural changes often go undetected in classical specifications, raising the risk of large drawdowns and poor capital allocation.
@@
For longer-term horizons, naive, shrinkage, and factor models are more common (Ledoit and Wolf,, 2004; De Nard et al.,, 2022). Naive approaches assume that recent covariance estimates strongly predict future covariances as they follow a Markov process. At the same time, shrinkage techniques blend the sample covariance matrix with structured targets to reduce noise, particularly when the cross-section of assets is large (Ledoit and Wolf,, 2022). Factor models can provide more insight but may not generalise well across diverse portfolios, especially when markets are affected by many different factors. These methods are typically stable but can underperform when major shifts occur.
@@
In recent years, there has been a surge in machine learning methods that can handle non-linearities and high dimensionality. Gu et al., (2020) and Kim et al., (2023) use techniques like random forests, support vector machines, and generative adversarial networks, while Zhang et al., (2024) applies graph neural networks to capture volatility spillovers for covariance matrix forecast. However, most of these solutions still emphasise short-term horizons.
@@
Attention mechanisms (Vaswani et al.,, 2017) offer a promising direction by weighting key observations more heavily. Some studies apply these techniques in finance but still focus on short-horizon tasks (Nazareth and Ramana Reddy,, 2023; Olorunnimbe and Viktor,, 2023). In principle, attention-based models could capture the slow-moving, cyclical factors and abrupt shifts that define medium- and long-horizon covariance forecasts, realising this potential. However, it requires integrating sophisticated data-driven architectures with the economic logic of excess returns.
@@
Recent efforts to extend these approaches to multi-asset portfolios have been limited. While many studies handle large cross-sections of stocks or focus on single-asset classes such as equities (Reis et al.,, 2024), fewer examine the combined dynamics across multiple asset classes in an integrated framework. However, institutional investors routinely diversify across equities, bonds, alternatives, and derivatives, underscoring the need for robust, multi-asset covariance forecasting models. In particular, capturing time-varying cross-asset correlations over medium-term horizons is a non-trivial challenge that remains under-explored in much of the existing literature.
@@
Two key research gaps remain. First, most existing studies focus on short- or long-term horizons, leaving a relative void in medium-term covariance forecasting. This gap is particularly salient for multi-asset portfolios, where cross-asset correlations over multi-week or multi-month windows can significantly influence asset allocation and risk management (De Nard et al.,, 2021). Second, there is limited empirical evidence on whether using raw returns versus excess returns leads to systematically different covariance forecasts, especially over these medium-term horizons. Economic theory suggests that risk premiums and the underlying drivers of asset co-movements could be more accurately captured through excess returns. However, most machine learning and econometric approaches mostly default to raw returns.
@@
Our study addresses these gaps by focusing on medium-term covariance forecasts in a multi-asset setting. Specifically, we propose a novel deep learning (DL) model that captures spatio-temporal correlations while remaining robust to structural changes. We benchmark against classical methods (naive, shrinkage, GARCH, and dimensional reduction) and demonstrate consistent gains in forecast accuracy under diverse market environments from 2017 to 2023. Our main contributions are as follows:
@@ 2 Methodology
To estimate ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}$, we compare naive and other benchmark models found in the literature alongside our proposed model. A description of each model is shown below.
2.1 Naïve ($NA$)
@@ 2.2 Naïve full sample($NA^{F}$)
2.3 Exponential Weighted Moving Average ($EWMA$)
@@
2.4 Ledoit and Wolf ($LW$)
The LW shrinkage (Ledoit and Wolf,, 2004) is a robust statistical model that combines the sample covariance matrix with a structured target matrix, improving estimation accuracy by reducing the impact of estimation error due to high dimensionality and limited sample size. The LW shrinkage estimator is given by:
@@ 2.4 Ledoit and Wolf ($LW$)
2.5 Ledoit and Wolf full sample ($LW^{F}$)
@@
2.6 Constant Conditional Correlations ($CCC$)
The $CCC$ model, introduced by Bollerslev, (1990), assumes that asset correlations remain constant over time while variances evolve dynamically according to a GARCH process based on a one-step forecast. However, since investment funds and risk managers typically rebalance portfolios over extended horizons, such as monthly or quarterly, a one-step forecast is insufficient for practical decision-making (De Nard et al.,, 2022).
To extend the model to a multi-step forecasting framework, we follow the methodology of Baillie and Bollerslev, (1992), defining the $F$-step ahead conditional covariance forecast as:
@@ 2.6 Constant Conditional Correlations ($CCC$)
2.7 Dynamic Conditional Correlations ($DCC$)
@@
2.8 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
One recent extension is the $DCC^{NL}$ model from De Nard et al., (2022). The covariance forecast in the $DCC^{NL}$ model follows the same structure as in Eq. (16), but now with a shrinkage-improved correlation matrix ($\overline{\mathbf{Q}}$). They apply the nonlinear shrinkage method of Ledoit and Wolf, (2020), which optimally shrinks the eigenvalues of the sample covariance matrix towards a better-conditioned target matrix. This procedure ensures that the estimated correlation matrix remains positive, definite, and well-conditioned, even in high-dimensional settings.
@@ 2.8 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
where $\mathbf{U}$ is the matrix of eigenvectors and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})$ is the diagonal matrix of eigenvalues. The nonlinear shrinkage transformation replaces $\mathbf{\Lambda}$ with a shrinkage-transformed version $\mathbf{\Lambda}^{*}$:
@@
2.9 Principal Component Analysis ($PCA$)
As noted by Ledoit and Wolf, (2003), empirical results show the usefulness of PCA in covariance matrix estimation. The PCA is computed as follows:
To ensure we capture most of the information, we retain only the first $k$ principal components that account for at least 95% of the total variance. This is mathematically expressed as:
$\sum_{i=1}^{k}\lambda_{i}\geq 0.95\sum_{i=1}^{N}\lambda_{i}$ (24)
The final forecasted covariance matrix is then approximated as:
2.10 Random Matrix Theory ($RMT$)
$RMT$ (Laloux et al.,, 2000) is built upon the $PCA$ where the upper ($\varsigma_{+}$) and lower ($\varsigma_{-}$) bound of the eigenvalues are calculated using the Marchenko-Pastur distribution. These bounds are calculated as:
This filtering process removes noise and retains only the significant components of the covariance matrix. The cleaned covariance matrix is reconstructed by multiplying the filtered eigenvalues with their corresponding eigenvectors:
@@ 2.11.3 Bidirectional LSTM (BiLSTM) Layer
The traditional BiLSTM outputs the last time step of the last layer. However, in our model, we use a similar approach to Liu and Guo, (2019) methodology, meaning all time steps from the last layer ($\mathbf{H}^{u}$) are used as an input to our following steps. To mitigate overfitting and enhance the model’s generalisation capabilities, a Dropout Layer is applied to the BiLSTM output. This operation randomly zeroes a fraction of the elements in $\mathbf{H}^{u}$ with a probability of 20%, effectively regularising the model by preventing reliance on specific neurons during training. During our testing phase, this feature will be skipped.
@@ 2.11.6 Shrinkage
While DL-based forecasts can learn intricate patterns, historical covariance may still convey valuable stability and continuity, particularly when the model is tested on data different from training. Therefore, our model’s final step is incorporating a linear shrinkage from historical values in our predictions.
@@ 3 Empirical Results
We conduct a rolling estimation procedure to simulate live conditions. Each new trading day $t$ in the test set triggers an online update of our DL model, incorporating all data up through day $t$. This design ensures that the forecasts mimic a real-world trading environment, where only up-to-date historical data is available at each decision point.
Table 2 reports the average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances between forecasted and realised covariance matrices for all models over the entire testing period (2021–2023) [^4] Bold and underlined numbers represent the lowest and the second lowest values, respectively.. GARCH models ($CCC$ and $DCC$) and $EWMA$ perform relatively well as they adapt faster to structural changes in the covariance matrices. However, our model ($CAB$) yields the lowest errors in both metrics, outperforming classical benchmarks. Dimensional reduction models perform inadequately as we only use 14 assets.
@@ 5 Economic Value in Portfolio Management
Using a rolling estimation window, we generate 11 portfolio strategies (one for each forecast model) plus an equally weighted (1/$N$) benchmark. After estimating $\widehat{\mathbf{\Sigma}}_{t}$, we compute the GMV weights and hold them until the next rebalancing date. We then record the ex-post average portfolio return for each model ($m$) as:
@@
Overall, $CAB$ tends to exhibit moderate turnover, balancing the need to adjust to changing correlations without incurring unnecessary trades. This stability benefits investors who must limit transaction costs or adhere to regulatory constraints on turnover. These results underscore that advanced covariance estimation can yield economically meaningful improvements in risk control, particularly over the medium-term horizons favoured by many institutional investors. The $CAB$ model stands out by combining strong variance reduction with stable rebalancing behaviour, balancing responsiveness to market changes and turnover efficiency.

### v2

@@ 1 Introduction
Covariance matrix forecast is central in modern finance. It influences portfolio construction, risk management, and asset pricing models. Markowitz, (1952) introduced mean-variance optimisation, highlighting how correlation among assets shapes the risk-return trade-off. Beyond portfolio choice, covariance forecast underpins factor-based models, hedging strategies, and even debates on market efficiency, where accurate correlation forecasts can drive risk control and potential outperformance.
@@
However, GARCH-type approaches struggle with large cross-sections, falling prey to the curse of dimensionality (Engle et al.,, 2019). Tools like principal component analysis and random matrix theory (Ledoit and Wolf,, 2003; Laloux et al.,, 2000) have been effective for large datasets. Recent advances, such as dynamic conditional correlation with nonlinear shrinkage (Engle et al.,, 2019), perform well overall, as per De Nard et al., (2024). Still, most of these methods focus on very short-term horizons, often just one step ahead or one day ahead, and rely on high-frequency data.
@@
Long-horizon forecasts matter for institutional investors, such as pension and investment funds. They rebalance portfolios over weeks or months. Therefore, a one-step forecast is insufficient for practical decision-making (De Nard et al.,, 2022). Furthermore, slow-moving trends, such as equity–bond correlation shifts, can be overlooked if one relies solely on high-frequency econometric models. Sandoval and Franca, (2012) highlights how regime shifts or significant structural changes often go undetected in classical specifications, raising the risk of large drawdowns and poor capital allocation.
@@
For longer-term horizons, naive and factor models are more common (Ledoit and Wolf,, 2004; De Nard et al.,, 2022). Factor models can provide insightful forecasts but may not generalise well across diverse portfolios, especially when different portfolios are affected by many different factors. These methods are typically stable but can underperform when major shifts occur.
Naive approaches assume that recent covariance estimates strongly predict future covariances as they follow a Markov process. However, these tend to be noisy when the number of assets tends to be close to the number of observations ($N\approx T$). For this purpose, shrinkage techniques reduce noise by blending the sample covariance matrix with structured targets, particularly when the cross-section of assets is large (Ledoit and Wolf,, 2022).
@@
While these methods rely on advanced statistical and econometric methods, we have seen a recent surge in machine learning methods that can also handle non-linearities and high dimensionality with more accurate forecast. Gu et al., (2020) and Kim et al., (2023) use techniques like random forests, support vector machines, and generative adversarial networks, while Zhang et al., (2024) applies graph neural networks to capture volatility spillovers for covariance matrix forecast. However, most of these solutions still emphasise short-term horizons.
@@
Attention mechanisms (Vaswani et al.,, 2017) offer a promising direction by weighting key observations more heavily. Some studies apply these techniques in finance but still focus on short-horizon tasks (Nazaret and Reddy,, 2023; Olorunnimbe and Viktor,, 2023). In principle, attention-based models could capture the slow-moving, cyclical factors and abrupt shifts that define medium- and long-horizon covariance forecasts, realising this potential. However, it requires integrating sophisticated data-driven architectures with the economic logic of excess returns.
@@
Recent efforts to extend these approaches to multi-asset portfolios have been limited. While many studies handle large cross-sections of stocks or focus on single-asset classes (Reis et al.,, 2024), fewer examine the combined dynamics across multiple asset classes in an integrated framework. However, institutional investors routinely diversify across equities, bonds, alternatives, and derivatives, underscoring the need for robust, multi-asset covariance forecasting models. In particular, capturing time-varying cross-asset correlations over medium-term horizons is a non-trivial challenge that remains under-explored in much of the existing literature.
@@
Two key research gaps remain. First, most existing studies focus on short- or long-term horizons, leaving a relative void in medium-term covariance forecasting. This gap is particularly salient for multi-asset portfolios, where cross-asset correlations over multi-week or multi-month windows can significantly influence asset allocation and risk management (De Nard et al.,, 2021). Second, there is limited empirical evidence on whether using raw returns versus excess returns leads to systematically different covariance forecasts, especially over these medium-term horizons. Economic theory suggests that underlying drivers of asset co-movements could be more accurately captured through excess returns (Fama and French,, 1993), i.e. by removing the risk-free component from returns. However, most machine learning and econometric approaches mostly default to raw returns.
@@
Our study addresses these gaps by focusing on medium-term covariance forecasts in a multi-asset setting. Specifically, we propose a novel deep learning (DL) model that captures spatio-temporal correlations while remaining robust to structural changes. We benchmark against classical methods and demonstrate consistent gains in forecast accuracy under diverse market environments from 2017 to 2023. Our main contributions are as follows:
@@ 2 Methodology
To estimate ${\widehat{\mathbf{\Sigma}}}_{t+1:t+F}$, we compare 10 benchmark models divided into three groups (naive, de-noising and GARCH models) alongside our proposed model. A description of each model is shown below.
2.1 Naïve Models
2.1.1 Naïve ($NA$)
@@ 2.1.2 Naïve full sample ($NA^{F}$)
2.1.3 Exponential Weighted Moving Average ($EWMA$)
@@
2.2 De-noising Models
2.2.1 Principal Component Analysis ($PCA$)
As noted by Ledoit and Wolf, (2003), empirical results show the usefulness of PCA in covariance matrix forecast. The PCA is computed as follows:
where $\mathbf{U}$ is the matrix of eigenvectors and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\lambda_{2},...,\lambda_{N})$ is the diagonal matrix of eigenvalues.
To ensure we capture most of the information, we retain only the first $k$ principal components that account for at least 95% of the total variance. This is expressed as:
$\sum_{i=1}^{k}\lambda_{i}\geq 0.95\sum_{i=1}^{N}\lambda_{i}$ (11)
The final forecasted covariance matrix is then approximated as:
2.2.2 Random Matrix Theory ($RMT$)
$RMT$ (Laloux et al.,, 2000) is built upon the $PCA$ where the upper ($\lambda_{+}$) and lower ($\lambda_{-}$) bound of the eigenvalues are calculated using the Marchenko-Pastur distribution. These bounds are calculated as:
This filtering process removes noise and retains only the significant components of the covariance matrix. The cleaned covariance matrix is reconstructed by multiplying the filtered eigenvalues ($\mathbf{\Lambda}^{\text{filtered}}$) with their corresponding eigenvectors:
2.2.3 Ledoit and Wolf ($LW$)
The LW shrinkage (Ledoit and Wolf,, 2004) is a robust statistical model that combines the sample covariance matrix with a structured target matrix, improving forecast accuracy by reducing the impact of forecast error due to high dimensionality and limited sample size. The LW shrinkage estimator is given by:
@@ 2.2.3 Ledoit and Wolf ($LW$)
2.2.4 Ledoit and Wolf full sample ($LW^{F}$)
@@
2.3 GARCH Models
2.3.1 Constant Conditional Correlations ($CCC$)
The $CCC$ model, introduced by Bollerslev, (1990), assumes that asset correlations remain constant over time while variances evolve dynamically according to a GARCH process based on a one-step forecast. To extend the model to a multi-step forecasting framework, we follow the methodology of Baillie and Bollerslev, (1992), defining the $F$-step ahead conditional covariance forecast as:
@@ 2.3.1 Constant Conditional Correlations ($CCC$)
2.3.2 Dynamic Conditional Correlations ($DCC$)
@@
2.3.3 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
One work that combines the multivariate GARCH with a nonlinear shrinkage is the model of Engle et al., (2019). The covariance forecast in the $DCC^{NL}$ model follows the same structure as in Eq. (22), but now with a shrinkage-improved correlation matrix ($\overline{\mathbf{Q}}$). The original model uses the numerical method of Ledoit and Wolf, (2015). However, we apply the nonlinear shrinkage method of Ledoit and Wolf, (2020) to speed up the computations, which optimally shrinks the eigenvalues of the sample covariance matrix towards a better-conditioned target matrix.
@@ 2.3.3 Nonlinear Shrinkage Dynamic Conditional Correlations ($DCC^{NL}$)
The nonlinear shrinkage transformation replaces $\mathbf{\Lambda}$ with a shrinkage-transformed version $\mathbf{\Lambda}^{*}$:
@@ 2.4.3 Bidirectional LSTM (BiLSTM) Layer
The traditional BiLSTM outputs the last time step of the last layer. However, in our model, we use a similar approach to Liu and Guo, (2019) methodology, meaning all time steps from the last layer ($\mathbf{H}^{u}$) are used as an input to our following steps. This procedure allows the attention mechanism to dynamically weigh the importance of each time step, enabling the model to focus on the most relevant temporal patterns when making predictions. By leveraging all time steps from the last layer ($\mathbf{H}^{u}$), rather than just the final output, the model can more effectively capture long-term dependencies and contextual relationships.
To mitigate overfitting and enhance the model’s generalisation capabilities, a Dropout Layer is applied to the BiLSTM output. This operation randomly zeroes a fraction of the elements in $\mathbf{H}^{u}$ with a probability of 20%, effectively regularising the model by preventing reliance on specific neurons during training. During our testing phase, this feature will be skipped.
@@ 2.4.6 Shrinkage
While DL-based forecasts can learn intricate patterns, which provide a more stable forecast, the noisy historical covariance may still convey valuable information about recent market conditions. Therefore, our model’s final step is incorporating a linear shrinkage from historical values in our predictions.
@@ 3 Empirical Results
We conduct a rolling forecast procedure to simulate live conditions. Each new trading day $t$ in the test set triggers an online update of our DL model, incorporating all data up through day $t$. This design ensures the forecasts mimic a real-world trading environment, where only up-to-date historical data is available at each decision point.
Table 2 reports the average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances between forecasted and realised covariance matrices for all models over the entire testing period (2021–2023) [^4] Bold and underlined numbers represent the lowest and the second lowest values, respectively.. GARCH models ($CCC$ and $DCC$) and $EWMA$ perform relatively well as they adapt faster to structural changes in the covariance matrices. Please note that although $DCC^{NL}$ is not more accurate than the traditional $DCC$, it provides more stable forecasts based on the eigenvalues. However, our model ($CAB$) yields the lowest errors in both metrics, outperforming classical benchmarks. Dimensional reduction models perform inadequately as we only use 14 assets.
@@ 4 Robustness analysis
4.1 Larger dataset
In this section, we expand our asset universe to 260 ETFs to evaluate whether our model continues to outperform established benchmarks in a high-dimensional setting. Appendix LABEL:tab:listofetfslarge provides a comprehensive list of the ETFs in a multi-asset portfolio. Table 4 presents the corresponding performance metrics.
Table 4: Average Euclidean ($\mathcal{L}^{E}$) and Frobenius ($\mathcal{L}^{F}$) distances (×$\times 10^{3}$) using 260 ETFs for $F=20$ and excess returns for different testing periods: Overall (1 January 2021 to 31 December 2023), Bull-1 (1 January 2021 to 2 January 2022), Bear (3 January 2022 to 12 June 2022), and Bull-2 (13 June 2022 to 31 December 2023).
Overall Bull-1 Bear Bull-2
$NA$ 11.546 16.484 8.677 12.442 20.458 29.257 10.814 15.390
$NA^{F}$ 10.792 15.399 7.661 11.014 20.742 29.532 9.927 14.130
$EWMA$ 10.544 15.054 7.518 10.802 18.410 26.292 10.219 14.543
$LW$ 10.719 15.304 8.095 11.619 19.933 28.458 9.740 13.867
$LW^{F}$ 13.460 19.170 12.098 17.270 19.374 27.609 12.622 17.947
$CCC$ 10.260 14.647 7.368 10.587 18.723 26.687 9.674 13.777
$DCC$ 10.280 14.671 7.796 11.182 18.076 25.777 9.622 13.703
$DCC^{NL}$ 11.047 15.747 8.597 12.301 18.687 26.634 10.414 14.814
$PCA$ 11.658 16.643 8.743 12.535 20.625 29.500 10.939 15.567
$RMT$ 11.652 16.636 8.738 12.529 20.618 29.490 10.934 15.561
$CAB$ 8.933 12.789 5.447 7.893 15.722 22.456 9.220 13.153
Our proposed approach ($CAB$) consistently outperforms the benchmark models, reflected by its lower average Euclidean and Frobenius distances across most periods. Notably, during the Overall, Bull-1, and Bull-2 periods, the post-hoc Nemenyi test confirms that the outperformance of CAB is statistically significant. In contrast, during the Bear period, CAB’s performance is statistically similar to that of the $DCC$ and $EWMA$ models.
@@ 5 Economic Value in Portfolio Management
Using a rolling forecast window, we generate 11 portfolio strategies (one for each forecast model) plus an equally weighted (1/$N$) benchmark. After estimating $\widehat{\mathbf{\Sigma}}_{t}$, we compute the GMV weights and hold them until the next rebalancing date. We then record the ex-post average portfolio return for each model ($m$) as:
@@
Table 7: Performance of the GMV portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio for the different market regimes. The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively.
Daily Weekly Monthly
$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
Bull-1 $1/N$ 0.009883 0.005169 0.009873 0.002335 0.009907 0.001103
$NA$ 0.000543 0.150112 0.000794 0.060372 0.000834 0.026885
$NA^{F}$ 0.000731 0.001378 0.000732 0.000607 0.000736 0.000327
$EWMA$ 0.000735 0.088541 0.000748 0.043884 0.000754 0.015613
$PCA$ 0.000850 0.236835 0.000827 0.090803 0.000795 0.020906
$RMT$ 0.000856 0.150601 0.000860 0.067402 0.000826 0.025859
$LW$ 0.001054 0.091979 0.001572 0.044883 0.001692 0.022283
$LW^{F}$ 0.000780 0.001228 0.000781 0.000610 0.000783 0.000349
$CCC$ 0.000620 0.119486 0.000681 0.053221 0.000728 0.013605
$DCC$ 0.000529 0.207262 0.000698 0.093558 0.000781 0.024324
$DCC^{NL}$ 0.000522 0.217432 0.000784 0.097809 0.000875 0.034673
$CAB$ 0.000634 0.086489 0.000656 0.055931 0.000707 0.023736
Bear $1/N$ 0.029939 0.007221 0.029940 0.003443 0.029793 0.001599
$NA$ 0.002207 0.092111 0.002682 0.032018 0.002959 0.018142
$NA^{F}$ 0.002891 0.001650 0.002892 0.000935 0.002906 0.000436
$EWMA$ 0.002908 0.079065 0.002843 0.036314 0.003120 0.019997
$PCA$ 0.002800 0.114256 0.002792 0.043637 0.002864 0.011111
$RMT$ 0.002983 0.094396 0.002896 0.037508 0.003153 0.023308
$LW$ 0.003118 0.095723 0.003970 0.041663 0.004754 0.021610
$LW^{F}$ 0.003196 0.002035 0.003197 0.001043 0.003201 0.000570
$CCC$ 0.002497 0.129487 0.002646 0.043843 0.002590 0.012451
$DCC$ 0.002355 0.067768 0.002639 0.015372 0.002636 0.005911
$DCC^{NL}$ 0.002269 0.083567 0.002652 0.024277 0.002730 0.010308
$CAB$ 0.002273 0.030506 0.002464 0.022686 0.002656 0.016836
Bull-2 $1/N$ 0.019967 0.005666 0.019956 0.002382 0.019933 0.001334
$NA$ 0.002823 0.118997 0.003526 0.053457 0.003667 0.029686
$NA^{F}$ 0.003398 0.001426 0.003400 0.000577 0.003406 0.000261
$EWMA$ 0.003432 0.073812 0.003433 0.034460 0.003630 0.016887
$PCA$ 0.003793 0.194898 0.003678 0.063899 0.003597 0.024379
$RMT$ 0.003668 0.118539 0.003720 0.055498 0.003762 0.029837
$LW$ 0.003512 0.102531 0.004274 0.048313 0.004245 0.020765
$LW^{F}$ 0.003874 0.002113 0.003879 0.001123 0.003889 0.000788
$CCC$ 0.003233 0.129874 0.003635 0.047812 0.003671 0.015276
$DCC$ 0.002947 0.071444 0.003294 0.026871 0.003364 0.009476
$DCC^{NL}$ 0.002924 0.085242 0.003318 0.034070 0.003365 0.012661
$CAB$ 0.003323 0.087821 0.003547 0.052596 0.003885 0.029487
@@
Overall, $CAB$ tends to exhibit moderate turnover, balancing the need to adjust to changing correlations without incurring unnecessary trades. This stability benefits investors who must limit transaction costs or adhere to regulatory constraints on turnover. These results underscore that advanced covariance forecast can yield economically meaningful improvements in risk control, particularly over the medium-term horizons favoured by many institutional investors. The $CAB$ model stands out by combining strong variance reduction with stable rebalancing behaviour, balancing responsiveness to market changes and turnover efficiency.